What are the quantum numbers for the electrons in the lithium

atom (Z = 3) when the atom is in its ground state?


1
Expert's answer
2021-05-17T19:22:51-0400

Lithium has two 1s-electrons and one 2s-electron (1s22s1).


n is equal to the number of the subshell

l = 0 ... (n - 1) (for s-electrons l = 0)

ml = -l ... +l

ms = 1/2 or -1/2

Therefore,


Quantum numbers for 1s electron (1):

n = 1

l = 0

ml = 0

ms = 1/2

Quantum numbers for 1s electron (2):

n = 1

l = 0

ml = 0

ms = -1/2

Quantum numbers for 2s electron:

n = 2

l = 0

ml = 0

ms = 1/2